OpenStudy (anonymous):

what is the slope of the line when the slope is 4/5 and the y-intercept is (0,-2)

OpenStudy (anonymous):

I think you just said the answer.

OpenStudy (anonymous):

if u mean the equation use this formula y-y1 =m(x-x1) m is the slope plug the numbers

OpenStudy (anonymous):

i believe that's what they want

OpenStudy (anonymous):

y+2=4/5(x) y=(4/5)x-2

OpenStudy (anonymous):

he the whole question says what is the slope-intercept equation of the line below, so the y intercept = (0,-2) and the slope= 4/5

OpenStudy (ranga):

slope-intercept equation is y = mx + b where m is the slope and b is the y intercept. Here slope m = 4/5 and y intercept b = -2 So y = 4/5x - 2 is the answer.

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

yw